OpenStudy (unknownrandom):

How would you do this? I can tell the answer is going to be 1. 4-√x=3

OpenStudy (unknownrandom):

I think I figured it out. subtract 4 from each side and then square root each?

OpenStudy (cybershadow):

yep or u just rearrange the equation like sqrt(x) = 4-3

OpenStudy (cybershadow):

that will leave sqrt(x) = 1

OpenStudy (cybershadow):

now square both sides to get x

OpenStudy (cybershadow):

that will give x = 1^2 =1

OpenStudy (unknownrandom):

Thanks for the help!

OpenStudy (cybershadow):

No worries!